# Assessment of the Earth orientation parameter accuracy from concurrent VLBI observations

### v1

@@ Abstract
We have assessed accuracy of estimates of Earth orientation parameters (EOP) determined from several very long baseline interferometry (VLBI) observing programs that ran concurrently at different networks. We consider that the root mean square of differences in EOP estimates derived from concurrent observations is a reliable measure of accuracy. We confirmed that formal errors based on the assumption that the noise in observables is uncorrelated are close to useless. We found no evidence that advanced scheduling strategies with special considerations regarding the ability to better solve for atmospheric path in zenith direction applied for 1-hr single-baseline sessions have any measurable impact on the accuracy of EOP estimates. From this, we conclude that there is a certain limit in our ability to solve for the atmospheric path delay using microwave observations themselves and a scheduling strategy is not the factor that impairs accuracy of EOP determination. We determined that EOP errors vary with season, being smaller in winter and greater in summer. We found that the EOP errors are scaled with an increase in duration of an observing session as a broken power law with the power of -0.3 at durations longer than 2–4 hours, which we explain as manifestation of the presence of correlations in the atmospheric noise.
@@ 1 Introduction
It was demonstrated in the 1980s that Very Long Baseline Interferometry (VLBI) outperforms traditional optical techniques for determination of irregularities in the Earth rotation by more than an order of magnitude [29, 43]. Regular VLBI observations dedicated to the determination of the Earth orientation parameters (EOPs) commenced on January 04, 1984 and are running since then without interruption. In order to provide a meaningful interpretation to a given parameter adjusted from measurements, its error should be characterized. It turned out to be a difficult task to assess the accuracy of the EOP from VLBI observations. VLBI does not directly measure Earth rotation. It measures voltage at receivers, which is the thermal noise with a tiny admixture of the noise from an observed extragalactic radio source. Lengthy records of voltage data — the dataset that we process for this study has over $3\cdot 10^{17}$ samples, undergo several stages of data analysis, including non-linear procedures. The last stage uses group delays as an input and estimates EOP using least squares. If we know precisely the covariance matrix of noise in group delays, we can easily arrive at dispersions of EOP estimates using the classical error propagation law. We do not know. As a coarse approximation, we form diagonal terms of the input covariance matrix from estimates of uncertainties of group delays setting off-diagonal terms to zero. It was known since 1980s [10, 24] that uncertainties of estimates of EOP and any other parameters derived that way are underestimated in a wide range from a factor of 1.2 to 10. These estimates of uncertainties are traditionally labeled as “formal” staying short from telling that these estimates of uncertainties are just wrong.
@@ 2 Formalism of the Earth orientation parameters
Since space is three-dimensional, a rotation of a rigid body is described by three parameters. The classical formalism of rotational kinematic was established by Leonhard Euler, and these rotation angles bear his name. Considering that we estimate a small perturbation of Euler angles with respect to an a priori model, we can linearize equations of motion and write the coordinate transformation of a vector $\vec{r}_{{}_{T}}$ from the crust-fixed terrestrial coordinate system to the vector in the inertial celestial coordinate system $\vec{r}_{{}_{C}}$ in a form
@@
where ^$\widehat{\mathstrut\cal M}_{a}(t)$ is the a priori matrix of the Earth rotation and $\vec{q}_{e}(t)$ is the vector of small residual Euler angles with components that we will label as $E_{1}$, $E_{2}$, $E_{3}$. Historically, polar motion and Universal Time were determined with two totally different techniques. Polar motion was determined from observations of zenith telescopes with a micrometer that was graduated with arcseconds. Therefore, a natural unit of polar motion was arcsecond. Universal Time was determined by measuring the time of passing a star through the meridian with respect to a local clock using the eye-hear method or, later, using photoelectric registration. Therefore, a natural unit of Universal Time was second. Neither method has been used for decades. Internally, partial derivatives of VLBI time delay with respect to components of the Euler rotation vector are written in radians. Therefore, natural units of EOP determined with VLBI are radians. Since we will compare $E_{1}$, $E_{2}$, and $E_{3}$, we found it is extremely inconvenient to use different units for components of the same vector and apply artificial scaling factors. Since we have to change unit scaling factors anyway, we set all of them to one and adhere SI unit for angles in our work, namely, radian. For a reference: 1 nrad = 0.206 mas = 13.8 $\mu$sec. Euler angles are related to vintage polar motion $X_{p}$, $Y_{p}$, and UT1 parameters as
$\begin{array}[]{lcl}E_{1}&=&\mu\,Y_{p}\vspace{0.5ex}\\ E_{2}&=&\mu\,X_{p}\vspace{0.5ex}\\ E_{3}&=&\kappa\,{\rm UT1},\\ \end{array}$ (5)
where $X_{p}$ and $Y_{p}$ are expressed in arcseconds, UT1 in seconds, and scaling factors $\mu=\pi/(180\cdot 60\cdot 60)\approx 4.848\cdot 10^{-6}$, $\kappa=-\pi/(12\cdot 60\cdot 60)\cdot(\Omega_{n}+\Omega_{p})/\Omega_{n}\approx-7.292115\cdot 10^{-5}$. Here $\Omega_{n}$ is the nominal Earth rotation angular velocity and $\Omega_{p}$ is the precession rate in right ascension.
Although the fact that space is three-dimensional seems firmly established, it is not uncommon to encounter a statement that there are five parameters that describe Earth rotation, for instance, in https://hpiers.obspm.fr/eop-pc/earthor/EOP.html. It was Herring et al. [11] who suggested estimating daily nutation offsets in the mid-1980s as a temporary measure because by that time the time series of VLBI experiments was too short to separate the principal nutation term with a period of 18.6 years, annual term, and precession. Nutation can be considered as a component of quasi-diurnal harmonic oscillations in the polar motion that occupies a certain part of the spectrum. Nowadays, the time span of high-quality VLBI observations is sufficient for the separation of nutation constituents, and we are in a position to return to a classical approach for estimation of the nutation term amplitudes in the frequency domain set by James Bradley in 1748 [see 28] as it has been demonstrated by Petrov [18] and now routinely used for common VLBI analysis [22]. The high-frequency spectrum of the Earth rotation exhibits a minor variability due to the free core nutation, and this variability causes estimates of the EOP spectrum to evolve with an increase in the time interval used for estimation. However, since the free-core nutation has a period of around 433 days, this variability is slow, and it is sufficient to re-evaluate the spectrum of high-frequency EOP every 2–5 months to capture this signal using a global dataset, including newly acquired data. The anharmonic constituent of EOP is much more variable and has to be estimated over a much shorter period of time, 0.5–5 days, using datasets that cover these intervals.
@@
In the context of our work, we consider harmonic variations in the Earth’s rotation as constant for a given dataset, but Euler angles as freely varying with time. We restrict the scope of our analysis to three Euler angles estimated over intervals of 0.5 to 24 hours long.
@@ 3 Observations
We used all 24-hr geodetic multi-baseline VLBI data since 1980 April 11 through 2025 March 03 in our analysis, in total 7818 experiments. In addition, we processed data from a number of 1-hr single baseline geodetic campaigns [^1] A small number of experiments of these campaigns had more than two stations.. We split these data into several subsets of target data, reference data, and remaining data labeled as background geodetic data. We process them as well, but we report in this study only results of EOP determination from target experiments.
@@
  • 148 24-hr experiments of regular quad-band programs VGOS-OPS, VGOS-TEST, and VGOS-RD, collectively called in this study vo, that ran in 2019.0–2024.0. In these experiments antennas of 12–13 m diameter with the slewing speed up to $12^{\circ}/s$ dubbed as VLBI Global Observing System [VGOS; 15].
    were used. Fast slewing speed greatly simplifies schedule optimization, in particular, allowing for an increased distribution of observations among azimuth and elevation angles. The use of four bands spread over the 7 GHz bandwidth allows determination of group delay with a reported precision of 3–5 ps. The network included 13 stations. It varied from 3 to 11 stations in a given experiment, with a median of 7 stations.
@@
  • 81 dual-band 24-hr experiments from program CORE-A, thereafter contracted to ca in 1997.1–2000.5 [38, 39, 37] at a global network of 13 stations. It varied from 4 to 6 stations in a given experiment with the median of 6 stations.
  • 235 dual-band 24-hr experiments from concurrent NEOS-RAPID, IVS-R1, IVS-R4 programs [40] have been running since 1993.4 till present and are concurrent with one of the above-mentioned programs. For brevity, we contract program names to na and r1r4. Two 24-hr experiments are considered concurrent if the middle epochs of these experiments are within 1 hour. The network included 32 stations. It varied from 4 to 15 stations in a given experiment, with the median of 8 stations.
@@
  • 3275 dual-band 1-hr experiments at stations Kk and Wz from IVS-INT-1 and IVS-INT-00 program in 1999.9–2025.2. Observations at this baseline run 3–4 days a week. We call this program KkWz.
  • 423 quad-band 1-hr experiments at stations K2 and Ws from VGOS-INT-A program in 2020.2–2025.2. Observations at this baseline run 3–4 days a week. We call this program K2Ws.
  • 133 dual-band 1-hr experiments at stations Ts and Wz from IVS-INT-2 program in 2002.8–2017.0. Observations at this baseline run 2–3 days a week till the antenna was decommissioned. We call this program TsWz.
@@
The experiments listed above are not the only concurrent geodetic observations. There were other concurrent programs, but they either covered a short interval of time, for instance, a 15-day long CONT-17 program [1], or had fewer experiments. Since differences in statistics of various programs are expected to differ at a level of tens of percents, a significant dataset needs to be analyzed in order to make an inference about whether the difference in statistics are significant or not. For instance, to establish that the standard deviations of two samples that differ at 20% are significant at a 0.05 level, the F-test requires a sample size of at least 118.
@@ 3.1 Scheduling S22 campaign
During scheduling VLBI Intensive sessions, several criteria are usually taken into account to be optimized [32]. These include: (1) maximizing the total number of observations, (2) obtaining observations at a range of elevation angles to enable accurate estimation of atmospheric path delays in zenith direction, (3) ensuring a balanced distribution of observations across multiple sources, and (4) specifically including sources at low declinations and near the cusps of the mutual sky visibility [33].
While generally speaking, observations at longer baselines are more sensitive for measuring Earth’s rotation angle $E_{3}$, they also face limitations due to a reduced common visibility of the sky [33]. The $\sim\!8400$ km baseline between stations Mg and Ws presents certain geometric advantages. At this distance, observations near the zenith at one station correspond to observations at very low elevation angles at the other station (see Figure 2). This geometry is particularly beneficial for solving for residual atmospheric path delay in zenith direction, which remains the primary source of error in VLBI measurements. To effectively separate atmospheric zenith path delay from other parameters, such as station height and clock biases, it is essential to observe at a wide range of elevation angles.
Figure 2: Mutual visibility between Ws (left) and Mg (right) stations. The elevation of the other station is color-coded. The black line marks the visible horizon [35].
The s22 sessions were specifically designed to investigate improved estimation of residual atmospheric path delay in zenith direction. Each session is divided into 1-hour blocks that alternate between two scheduling strategies. Strategy A follows the conventional scheduling approach utilized in the VieSched++ scheduling software [31] and serves as a baseline reference. By contrast, strategy B is tailored to maximize the variability in elevation angles, promoting frequent alternation between high- and low-elevation observations to improve sensitivity to zenith atmospheric path delay parameters.
To implement Strategy B, sources at high elevation for each station are identified in advance. Scans are then scheduled according to the following repeating pattern: (1) a source at high elevation for Mg (corresponds to low elevation for Ws), (2) any source, (3) a source at high elevation for Ws (corresponds to low elevation for Mg), (4) any source.
The interleaved “any source” steps (2 and 4) help to maintain a balanced distribution of observations across different sources. In general, a minimum interval of 20 minutes is enforced between consecutive observations of the same source. However, due to the limited availability of high-elevation sources at each station, this restriction is relaxed to 10 minutes for those specific sources in strategy B. Additionally, to ensure adequate coverage of low-declination regions, a low-declination source is explicitly scheduled every 15 minutes in both strategies.
@@ 4 Data analysis
The first active signal amplification of the voltage at a receiver at radio frequency bands is made by the cryogenically cooled radio telescope’s front-end system. These radio frequency voltage signals of interest are typically down-converted to lower radio frequency (RF) bands for further handling. The data acquisition terminal typically comprises a digitizer for RF bands as a backend located in the operations room, which converts the analog voltage signal into a time and amplitude discrete digital representation of the frequency band-limited analog signal. These digital converter units extract a determined number of individual intermediate frequencies (IF) from each digitized RF-frequency band and convert them into the baseband channels. The individual channels are grouped as a set of channels together, parameterized by start frequency, bandwidth, and amplitude resolution. They are digitally formatted, time-tagged, and then recorded on a data storage for further processing. For example, a quad-band setup has 32 channels with 32 MHz channel bandwidth and 2-bit amplitude resolution for one linear polarization of the received radio wave. Thus, for both linear polarizations, horizontal and vertical, 64 streams of formatted data are recorded at a rate of 8 Gigisamples/s. All signal operations, from radio frequency conversion downstream to individual digital baseband channels, are coherently driven by the reference frequency of the Hydrogen maser and inherit its frequency stability.
Streams of raw voltages formatted samples are processed with the DiFX software correlator [4, 5]. It reads raw voltage samples, unpacks them from 2-bit amplitude sampling to a floating-point representation, shifts the data streams according to a coarse path delay model to some arbitrary reference points, performs Fourier transform, computes complex auto- and cross-spectra, averages them with time, and then writes the results that are called visibilities. In addition, time series with amplitude and phase of the phase calibration tones are also written into individual files for each scan and station. Results of this correlation processing are called Level 1A data [^2] See https://github.com/nasa/sgdass/blob/main/pima/doc/swin_format.pdf for specifications..
@@
$C(\tau_{p},\tau_{g},\dot{\tau}_{p})e^{-i2\pi\,f_{0}\tau_{p}}=$ $\displaystyle\sum_{k}\sum_{j}c_{kj}\times e^{i2\pi(f_{0}\dot{\tau}_{p}(t_{k}-t_{0})\;+\;(f_{j}-f_{0})\tau_{g})}.$ (6)
In a case of quad-band data, the fringe fitting process estimates phase, ionosphere-free group delay, and the contribution of the ionosphere in a single fit. In a case of dual-band data, we process visibilities from each band separately, and then form a linear combination of group delays with coefficients selected in such a way that the frequency-dependent contribution of path delay in the ionosphere is canceled [see, for instance, 21].
Fourfit computes the uncertainties of group delays based on the achieved signal to noise ratios (SNR) assuming individual samples of visibilities are independent and uncorrelated with time and frequency using the equations 6.45 and A12.24 presented in Thompson et al. [41]:
@@
where $\Delta f_{\rm rms}$ is the root mean square of IF within the band.
The ionosphere-free group delay or a combination of group delays is the basis of further data analysis.

### v2

@@ Abstract
We have assessed accuracy of estimates of Earth orientation parameters (EOP) determined from several very long baseline interferometry (VLBI) observing programs that ran concurrently at different networks. We consider that the root mean square of differences in EOP estimates derived from concurrent observations is a reliable measure of accuracy. We confirmed that formal errors based on the assumption that the noise in observables is uncorrelated   have a limited use. We found no evidence that advanced scheduling strategies with special considerations regarding the ability to better solve for atmospheric path in zenith direction applied for 1-hr single-baseline sessions have any measurable impact on the accuracy of EOP estimates. From this, we conclude that there is a certain limit in our ability to solve for the atmospheric path delay using microwave observations themselves and a scheduling strategy is not the factor that impairs accuracy of EOP determination. We determined that EOP errors vary with season, being smaller in winter and greater in summer. We got the quantitative estimate of the impact of unmodeled source structure on EOP estimates and we found that the seasonal extra variance is one order of magnitude greater than the impact of source structure. We   established that the EOP errors are scaled with an increase in duration of an observing session as a broken power law with the power of -0.3 at durations longer than 2–4 hours, which we explain as a manifestation of the presence of correlations in the atmospheric noise.
@@ 1 Introduction
It was demonstrated in the 1980s that Very Long Baseline Interferometry (VLBI) outperforms traditional optical techniques for determination of irregularities in the Earth rotation by more than an order of magnitude [r:merit82, r:merit00]. Regular VLBI observations dedicated to the determination of the Earth orientation parameters (EOPs) commenced on January 04, 1984 and are running since then without interruption. In order to provide a meaningful interpretation to a given parameter adjusted from measurements, its error should be characterized. It turned out to be a difficult task to assess the accuracy of the EOP from VLBI observations. VLBI does not directly measure Earth rotation. It measures voltage at receivers, which is the thermal noise with a tiny admixture of the noise from an observed extragalactic radio source. Lengthy records of voltage data — the dataset that we process for this study has over $3\cdot 10^{17}$ samples, undergo several stages of data analysis, including non-linear procedures. The last stage uses group delays as an input and estimates EOP using least squares. If we know precisely the covariance matrix of noise in group delays, we can easily arrive at dispersions of EOP estimates using the classical error propagation law. We do not know. As a coarse approximation, we form diagonal terms of the input covariance matrix from estimates of uncertainties of group delays setting off-diagonal terms to zero. It was known since 1980s [r:her83, r:ray91] that uncertainties of estimates of EOP and any other parameters derived that way are underestimated in a wide range from a factor of 1.2 to 10. These estimates of uncertainties are traditionally labeled as “formal” staying short from telling that these estimates of uncertainties are just wrong.
@@ 2 Formalism of the Earth orientation parameters
Since space is three-dimensional, a rotation of a rigid body is described by three independent parameters.   We follow the classical approach for describing rotation of a rigid body established in the XVIII century and that is introduced in modern physics textbooks [see, for instance, r:Goldstein_book, r:Taylor_book, r:Landau_Lifshitz_book]. In the Cartesian coordinate system a transformation from the inertial space to the body-fixed coordinate system is described by three Euler angles. Considering that we estimate a small perturbation of Euler angles with respect to an a priori model, we can linearize equations of motion and write the coordinate transformation of a vector $\vec{r}_{{}_{T}}$ from the crust-fixed terrestrial coordinate system to the vector in the inertial celestial coordinate system $\vec{r}_{{}_{C}}$ in a form
@@
where $\widehat{\mathstrut\cal M}_{a}(t)$ is the a priori matrix of the Earth rotation and $\vec{q}_{e}(t)$ is the   right-handed 3-vector in the Cartesian orthogonal basis of small residual Euler angles with components that we will label as $E_{1}$, $E_{2}$, $E_{3}$. Since this approach is based in classical physics textbooks, we call it classical. Historically, polar motion and Universal Time were determined with two totally different techniques. Polar motion was determined from observations of zenith telescopes with a micrometer that was graduated with arcseconds. Therefore,   an instrumental unit of polar motion was arcsecond. Universal Time was determined by measuring the time of passing a star through the meridian with respect to a local clock using the eye-hear method or, later, using photoelectric registration. Therefore,   an instrumental unit of Universal Time was second. Neither method has been used for decades. Internally, partial derivatives of VLBI time delay with respect to components of the Euler rotation vector are written and coded in radians. Therefore, natural units of EOP determined with VLBI are radians. We use radians in our work because this the SI standard unit for angle and because it is essential for our work to use the same units for all three components of the Earth rotation for their intercomparison.
The Earth orientation parameters in the notation that is currently used by the International Earth Rotation Service (IERS) can be converted to the classical Euler angles notation as
[…]
The vector of perturbations of Euler angles $\vec{q}_{e}$ is a smooth function of time. We can approximate it using simple analytical functions that depend on some numbers collectively known as Earth Orientation Parameters. A search of optimal parameterization is an important problem of geodesy [see, for instance, r:biz23]. Evaluation of harmonic variations from observations poses a certain difficulty. Historically, for 250 years amplitudes of the spectrum of harmonic variations of $E_{1},E_{2}$ in the retrograde diurnal band called nutations were estimated from observations. It was r:her86a who suggested estimating daily nutation offsets in the mid-1980s as a temporary measure to overcome degeneracy because at that time the time series of VLBI experiments was too short to separate the principal nutation term with a period of 18.6 years, the annual term, and precession rate. This approach is widely used and became rather common [r:iers2010]. Nowadays, the time span of high-quality VLBI observations is sufficient for the separation of nutation constituents, and we are in a position to return to a classical approach for estimation of the nutation term amplitudes in the frequency domain set by James Bradley in 1748 [see r:bradley] as it has been demonstrated by r:erm and nowadays routinely used for common VLBI analysis [see, for instance, r:jmg96, r:ibnu20, r:rfc1]. The detailed discussion of advantages and disadvantages of the parameterization of harmonic variations in the Earth rotation in a form of nutation daily offsets or in a form of complex amplitudes of the spectra. The transformation between two parameterizations can be found in r:erm and a brief summary is given in appendix B.
@@
In the context of our work, we consider harmonic variations in the Earth’s rotation that we solve for in our analysis as constant for a given dataset, but residual anharmonic Euler angles as freely varying with time. We restrict the scope of our analysis to three Euler angles estimated over intervals of 0.5 to 24 hours long.
@@ 3 Observations
We used all 24-hr geodetic multi-baseline VLBI data since 1980 April 11 through 2025 March 03 in our analysis, in total 7818 experiments.
In addition, we processed data from a number of 1-hr single baseline geodetic campaigns [^1] A small number of experiments of these campaigns had more than two stations.. We split these data into several subsets of target data, reference data, and remaining data labeled as background geodetic data. We process them as well, but we report in this study only results of EOP determination from target experiments.
@@
  • 148 24-hr experiments of regular quad-band programs VGOS-OPS, VGOS-TEST, and VGOS-RD, collectively called in this study vo, that ran in 2019.0–2024.0. In these experiments antennas of 12–13 m diameter with the slewing speed up to $12^{\circ}/s$ dubbed as VLBI Global Observing System [VGOS; r:nie18] were used. Fast slewing speed greatly simplifies schedule optimization, in particular, allowing for an increased distribution of observations among azimuth and elevation angles. The use of four bands spread over the 7 GHz bandwidth allows determination of group delay with a reported precision of 3–5 ps. The network included 13 stations. It varied from 3 to 11 stations in a given experiment, with a median of 7 stations.
@@
  • 81 dual-band 24-hr experiments from program CORE-A, thereafter contracted to ca in 1997.1–2000.5 [r:core99, r:core01, r:core02] at a global network of 13 stations. It varied from 4 to 6 stations in a given experiment with the median of 6 stations.
  • 235 dual-band 24-hr experiments from concurrent NEOS-RAPID, IVS-R1, IVS-R4 programs [r:r1r4] have been running since 1993.4 till present and are concurrent with one of the above-mentioned programs. For brevity, we contract program names to na and r1r4. Two 24-hr experiments are considered concurrent if the middle epochs of these experiments are within 1 hour. The network included 32 stations. It varied from 4 to 15 stations in a given experiment, with the median of 8 stations.
@@
  • 3275 dual-band 1-hr experiments at stations Kk and Wz from IVS-INT-1 and IVS-INT-00 program in 1999.9–2025.2. Observations at this baseline run   3–5 days a week. We call this program KkWz.
  • 423 quad-band 1-hr experiments at stations K2 and Ws from VGOS-INT-A program in 2020.2–2025.2. Observations at this baseline run   3–5 days a week. We call this program K2Ws.
  • 133 dual-band 1-hr experiments at stations Ts and Wz from IVS-INT-2 program in 2002.8–2017.0. Observations at this baseline run 2–3 days a week till the Ts antenna was decommissioned. We call this program TsWz. Since 2023 observations between Wz and station Is that is located close to Ts resumed. We call that program IsWz, and we included 65 such experiments.
@@
  • 58 dual-band 1-hr experiments at stations Hn and Wz from USNO-INT-P program.
@@
The experiments listed above are not the only concurrent geodetic observations. There were other concurrent programs, but they either covered a short interval of time, for instance, a 15-day long CONT-17 program [r:cont17_org], or had fewer experiments. Since differences in statistics of various programs are expected to differ at a level of tens of percents, a significant dataset needs to be analyzed in order to make an inference about whether the difference in statistics are significant or not. For instance, to establish that the standard deviations of two samples that differ at 20% are significant at a 0.05 level, the F-test [r:fisher25] requires a sample size of at least 118.
@@ 3.1 Scheduling S22 campaign
During scheduling VLBI Intensive sessions, several criteria are usually taken into account to be optimized [r:vips++a]. These include: (1) maximizing the total number of observations, (2) obtaining observations at a range of elevation angles to enable accurate estimation of atmospheric path delays in zenith direction, (3) ensuring a balanced distribution of observations across multiple sources, and (4) specifically including sources at low declinations and near the cusps of the mutual sky visibility [r:scha21].
We used two strategies for scheduling in each experiment: strategy A for segments with odd indices and strategy B for segments with even indices. They are described in detail in appendix A.
@@ 4 Data analysis
Conventional geodetic instruments like theodolites or laser range meters perform direct observations, and assessment of errors of such individual observation is intuitively clear. VLBI techniques does not measure Earth rotation directly, and the error propagation through a very complicated pipeline is rather obscure.
The first active signal amplification of the voltage at a receiver at radio frequency bands is made by the cryogenically cooled radio telescope’s front-end system. These radio frequency voltage signals of interest are typically down-converted to lower radio frequency (RF) bands for further handling. The data acquisition terminal typically comprises a 2-bit digitizer for RF bands as a backend located in the operations room, which converts the analog voltage signal into a time and amplitude discrete digital representation of the frequency band-limited analog signal. These digital converter units governed by the Hydrogen maser extract a determined number of individual intermediate frequencies (IF) from each digitized RF-frequency band and convert them into the baseband   steams that are written alongside with time stamps.
Contribution of emission from an observed radio source and thermal emission of the receiver itself are very close to the white noise, and the root mean square (rms) of the latter noise is typically three to five orders of magnitudes greater than the rms of the emission from the radio source.
Streams of raw voltages formatted samples are processed with the DiFX software correlator [r:difx1, r:difx2]. It reads raw voltage samples, unpacks them from 2-bit amplitude sampling to a floating-point representation, shifts the data streams according to a coarse path delay model to some arbitrary reference points, performs Fourier transform, computes complex auto- and cross-spectra, averages them with time, and then writes the results that are called visibilities. Results of this correlation processing are called Level 1A data [^2] See https://github.com/nasa/sgdass/blob/main/pima/doc/swin_format.pdf for specifications..
@@
$\hskip-10.00002ptC(\tau_{p},\tau_{g},\dot{\tau}_{p})e^{-i2\pi\,f_{0}\tau_{p}}=$ $\displaystyle\sum_{k}\sum_{j}c_{kj}\times e^{i2\pi(f_{0}\dot{\tau}_{p}(t_{k}-t_{0})\;+\;(f_{j}-f_{0})\tau_{g})}.$ (6)
Since thermal noise from two datastreams is independent, but the common emission from an observed source is the same, averaging reduces the impact of the thermal noise to $C(\tau_{p},\tau_{g},\dot{\tau}_{p})$.
In a case of quad-band data, the fringe fitting process estimates phase, ionosphere-free group delay, and the contribution of the ionosphere in a single fit. In a case of dual-band data, we process visibilities from each band separately, and then form a linear combination of group delays with coefficients selected in such a way that the frequency-dependent contribution of path delay in the ionosphere is canceled [see, for instance, r:sba].
Fourfit computes the uncertainties of group delays based on the achieved signal to noise ratio (SNR) of $C(\tau_{p},\tau_{g},\dot{\tau}_{p})$ amplitudes assuming individual samples of visibilities are independent and uncorrelated with time and frequency using the equations 6.45 and A12.24 presented in r:tms:
@@
where $\Delta f_{\rm rms}$ is the root mean square of IF within the band. We have to stress that $\sigma_{\rm gr}$ in expression 7 is based on the assumption that the noise process affects phase and amplitude of $C$ equally. This assumption is valid for specific noise processes. For instance, the delay uncertainly inversely depends on the rms of receiver thermal noise. However, processes like delay in the propagation media and signal chain affect amplitude in a much lesser extent than phase. If phase variations during an observation are $\Delta\phi\gg 2\pi$, then their impact on group delay delay is proportional to $\Delta\phi^{-1}$, but the impact on amplitude is proportional to $\Delta\phi^{-2}$. Therefore, group delay uncertainty computed using expression 7 does not fully account the contribution of the phase noise, unless it causes phase variations during an observations comparable to a phase turn. This is one of the factors that eventually leads to an underestimation of EOP estimate errors.
The ionosphere-free group delay or a combination of group delays   and their uncertainties are the basis of further data analysis.